# The Brownian loop measure on Riemann surfaces and applications to length spectra

### v2

@@ Abstract
We prove a simple identity relating the length spectrum of a Riemann surface to that of the same surface with an arbitrary number of additional cusps. Our proof uses the Brownian loop measure introduced by Lawler and Werner. In particular, we express the total mass of Brownian loops in a fixed free homotopy class on any Riemann surface in terms of the length of the geodesic representative for the complete constant curvature metric. This expression also allows us to write the electrical thickness of a compact set in $\mathbb{C}$ separating 00 and $\infty$, or the Velling–Kirillov Kähler potential, in terms of the Brownian loop measure and the zeta-regularized determinant of Laplacian as a renormalization of the Brownian loop measure with respect to the length spectrum.
@@ 1 Introduction
Let $(X,g)$ be a Riemannian surface without boundary (or we replace $X$ with boundary by $X\setminus\partial X$). From the uniformization theorem, if $X$ is not one of ${}^\widehat{\mathbb{C}}=\mathbb{C}\cup\{\infty\}$, $\mathbb{C}$, $\mathbb{C}\setminus\{0\}$, and tori, then $X$ admits a unique complete hyperbolic metric that is conformal to $g$. In particular, $X$ may have cusps and funnels and may be of infinite type. We say a free homotopy class of oriented closed curves on $X$ is essential if it is not a trivial homotopy class nor a class homotopic to a cusp. Every such homotopy class of oriented closed curves on $X$ contains a unique hyperbolic geodesic $\gamma$. We define the iteration number $m(\gamma)\in\mathbb{Z}_{>0}$ if $\gamma$ is $m(\gamma)$-time iteration of a primitive closed geodesic.
@@ 1.2 Explicit computations of Brownian loop measure
Another immediate consequence is the following relation between the electrical thickness [22] and Brownian loop measure. Let $K\subset\mathbb{C}\setminus\{0\}$ be a connected compact set separating 00 and $\infty$. Let $\Omega$ and $\Omega^{*}$ be respectively the connected component of $\mathbb{C}\setminus K$ containing 00 and $\infty$. Let $f$ be a conformal map $\mathbb{D}\to\Omega$ fixing 00 and $h$ a conformal map $\mathbb{D}^{*}\to\Omega^{*}$ fixing $\infty$, where $\mathbb{D}=\{z\in\widehat{\mathbb{C}}\,|\,|z|<1\}$ and $\mathbb{D}^{*}=\{z\in\widehat{\mathbb{C}}\,|\,|z|>1\}$. The electrical thickness of $K$ is defined as
@@
Let $K\subset\mathbb{C}\setminus\{0\}$ be a connected compact set separating 00 and $\infty$. For all $m\in\mathbb{Z}\setminus\{0\}$, we have
@@
$\mu_{T}\{\text{loops homotopic to }\gamma^{m}\text{ but do not hit }\gamma\}=\frac{\operatorname{Area}(T)}{\pi\ell_{\gamma}(T)^{2}m^{2}}-\frac{1}{|m|}\frac{1}{\exp(\frac{\pi\ell_{\gamma}(T)^{2}|m|}{\operatorname{Area}(T)})-1},$
@@ 1.3 Zeta-regularized determinant of Laplacian
We show another way to express $-\log\mathrm{det}_{\zeta}\Delta_{g}$ on a closed hyperbolic surface $X$ as a renormalization of the Brownian loop measure. The following result is a rewriting of the Selberg trace formula, which builds a connection between the Laplacian spectrum and the length spectrum, and hence is helpful to translate $\log\mathrm{det}_{\zeta}\Delta$ into the language of Brownian loop measure. Denote $\mathcal{G}(X)$ the set of all oriented closed geodesics on $X$, and $\mathcal{P}(X)\subset\mathcal{G}(X)$ the subset of primitive ones. Let $N_{X}(L):=\#\{\gamma\in\mathcal{P}(X)\mid\ell(\gamma)\leq L\}$.
@@
We note that the integral $\int_{L=0}^{\infty}\frac{1}{\mathrm{e}^{L}-1}\ \mathrm{d}N_{X}(L)$ is formally the total mass of Brownian loops homotopic to a primitive geodesic by Lemma 1.1, although infinite. This result expresses $\log\mathrm{det}_{\zeta}\Delta$ of a closed hyperbolic surface $X$ only in terms of its genus (since Area$(X)=2\pi(2\operatorname{genus}(X)-2)$ by the Gauss–Bonnet theorem) and a rather simple function of the length spectrum of $X$. Moreover, the normalizing integral with respect to $\widetilde{\mathop{\rm Li}}(\mathrm{e}^{L})$ is universal (does not depend on $X$).
@@ 3 Brownian loop measure in a homotopy class
Since the Brownian loop measure is conformally invariant, choosing a complete metric with constant curvature on $X$ allows us to explicitly compute the total mass. In particular, $X$ is stochastically complete (i.e., Brownian motion is defined for infinite time, see Appendix A). If $X$ is not hyperbolic, then $X$ is conformal to one of ^,,$\widehat{\mathbb{C}},\mathbb{C},\mathbb{C}\setminus\{0\}$ or flat tori. On $\widehat{\mathbb{C}},\mathbb{C}$ and $\mathbb{C}\setminus\{0\}$, every homotopy class of closed curves is either homotopically trivial or homotopic to a puncture point, and hence the mass of Brownian loops is infinite. So, we will only consider complete hyperbolic surfaces and flat tori.
@@ 3.1 On hyperbolic surfaces
Let $X$ be a Riemann surface without boundary endowed with the unique complete hyperbolic metric. Then it is isometric to a quotient of the upper half plane $\𝑋\ΓX=\Gamma\backslash\mathbb{H}$, where $\Gamma$ is a Fuchsian group (i.e., a discrete subgroup of ${\rm PSL}(2,\mathbb{R})$ which acts on $\mathbb{H}$ by $\big{(}\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big{)}:z\mapsto\frac{az+b}{cz+d}$) without elliptic elements. There exists a unique closed geodesic in every essential (i.e., non-trivial and not homotopic to a cusp) free homotopy class. Let $\gamma$ be an oriented closed geodesic on $X$, and denote $\ell_{\gamma}(X)=\ell(\gamma)$ be the length of it. We write
$\mathcal{C}_{X}(\gamma):=\{\eta\,|\,\eta\text{ is a closed curve in $X$ free homotopic to }\gamma\}.$
@@
where $[\gamma]$ is the conjugacy class of $\gamma$ in $\Gamma$ and $z\in\mathbb{H}$ is any lift of $x$.
@@
Let $\𝑋\ΓX=\Gamma\backslash\mathbb{H}$ be a complete hyperbolic surface without boundary. Let $\gamma$ be an oriented closed geodesic on $X$ of length $\ell(\gamma)$ and iteration number $m(\gamma)$. We have
@@
Fix an element $h_{0}\in[\gamma]$ and then $[\gamma]=\{h_{1}^{-1}h_{0}h_{1}\mid h_{1}\in\Gamma\}$. Two elements $h_{1}^{-1}h_{0}h_{1}=h_{2}^{-1}h_{0}h_{2}$ if and only if $h_{2}h_{1}^{-1}h_{0}=h_{0}h_{2}h_{1}^{-1}$, that is, $h_{2}h_{1}^{-1}$ is in the centralizer of $h_{0}$ in $\Gamma$. Since $\Gamma$ is a discrete subgroup of ${\rm PSL}(2,\mathbb{R})$, the centralizer of $h_{0}\in\Gamma$ is just the set $\left\langle\delta\right\rangle:=\{\delta^{k}\mid k\in\mathbb{Z}\}$ where $\delta$ is primitive and $h_{0}=\delta^{m(\gamma)}$. Then
@@
where for each class $[h_{1}]\in\left\langle\delta\right\rangle\backslash\Gamma$ we may choose any representation element.
@@
$U(u)$ ::=:= $\int_{\mathop{\rm arccosh}(u^{2})}^{\infty}\frac{r\mathrm{e}^{-r^{2}/4t}}{\sqrt{\cosh{r}-u^{2}}}\,\mathrm{d}r$ (3.3)
@@
[…]
@@ 4.2 On surfaces with boundary
As we mentioned at the beginning of Section 3, we may also use Lemma 3.2 to compute the Brownian loop measure on a surface $X$ with a non-empty boundary (where the Brownian loops are killed while hitting the boundary). Let $X^{\prime}=X\setminus\partial X$. Since $X^{\prime}$ cannot be one of ^,,$\widehat{\mathbb{C}},\mathbb{C},\mathbb{C}\setminus\{0\}$ or flat torus, it admits a unique complete hyperbolic metric. Then, by Theorem 2.2 and Lemma 3.2, the Brownian loop measure can be computed on $X^{\prime}$ with the complete hyperbolic metric:
@@
Let $K\subset\mathbb{D}$ be a closed subset such that $D=\mathbb{D}\setminus K$ is simply connected and contains 00. Let $\psi:D\to\mathbb{D}$ be a conformal map fixing 00. The total mass of Brownian loop in $\mathbb{D}$ intersecting $K$ with winding number $m$ around 00 equals $\log|\psi^{\prime}(0)|/(2\pi^{2}m^{2})$.
@@
Since $\psi$ is conformal, for all $\varepsilon>0$, there exists $r_{0}$ large enough, such that for all $r>r_{0}$, the curve $\psi(\mathrm{e}^{-r}\mathbb{S}^{1})$ is contained in the annulus $\{z\in\mathbb{C}\,|\,|\psi^{\prime}(0)|\mathrm{e}^{-r-\varepsilon}<|z|<|\psi^{\prime}(0)|\mathrm{e}^{-r+\varepsilon}\}$. See Figure 2 for an illustration. As $\mathbb{A}_{r^{\prime}}$ is conformally equivalent to the annulus bounded by $\mathbb{S}^{1}$ and $\psi(\mathrm{e}^{-r}\mathbb{S}^{1})$, we have from the monotonicity of the modulus of annuli for inclusion that $-\log|\psi^{\prime}(0)|+r-\varepsilon<r^{\prime}<-\log|\psi^{\prime}(0)|+r+\varepsilon$. This implies $\lim_{r\to\infty}r-r^{\prime}=\log|\psi^{\prime}(0)|$.
@@
Let $K\subset\mathbb{C}\setminus\{0\}$ be a compact set separating 00 and $\infty$. Let $\Omega$ and $\Omega^{*}$ be respectively the connected component of $\mathbb{C}\setminus K$ containing 00 and $\infty$. Let $f$ be a conformal map $\mathbb{D}\to\Omega$ fixing 00 and $h$ a conformal map $\mathbb{D}^{*}\to\Omega^{*}$ fixing $\infty$. The electrical thickness [22] of $K$ (or the Velling–Kirillov Kähler potential in the context of universal Teichmüller theory [50]) is defined as
@@
Let $K\subset\mathbb{C}\setminus\{0\}$ be a connected compact set separating 00 and $\infty$. For all $m\in\mathbb{Z}\setminus\{0\}$, we have
@@
We write $\iota:z\mapsto 1/z$ and let $\tilde{h}=\iota\circ h\circ\iota$ which maps $\mathbb{D}$ conformally onto $\iota(\Omega^{*})$ while fixing 00. Then we have $|h^{\prime}(\infty)|=1/|\tilde{h}^{\prime}(0)|$ and $\theta(K)=-\log|\tilde{h}^{\prime}(0)|-\log|f^{\prime}(0)|$. The left-hand side of (4.6) equals
@@
Since $X$ is not conformally equivalent to one of ^,,$\widehat{\mathbb{C}},\mathbb{C},\mathbb{C}\setminus\{0\}$ or closed tori, it admits a complete hyperbolic metric. We have $\𝑋\ΓX=\Gamma\backslash\mathbb{H}$ where $\Gamma$ is a Fuchsian group without elliptic elements. Let $\mathcal{P}(X)$ be the set of all primitive, oriented, closed geodesics on $X$. By Lemma 3.2, the total mass of Brownian loops of all essential homotopy classes is
@@
where $\delta>0$ is the Hausdorff dimension of the limit set of $\Gamma$. If $\delta>\frac{1}{2}$, then $\delta(1-\delta)$ is the smallest eigenvalue of $-\Delta$ (acting on $L^{2}(X)$) on $X$ [40, 47, 48, 49]. In our case, $X$ has infinite area and hence constant valued functions are not $L^{2}$-eigenfunctions. So it has no zero eigenvalue (since the eigenfunction of zero eigenvalue is of constant value), and $\delta$ is strictly smaller than $1$. Then, the total mass of Brownian loops (4.7) is finite. ∎
@@ 4.3 Zeta-regularized determinant of Laplacian
Let us first recall briefly the definition of the zeta-regularized determinant. The spectrum of the positive Laplace–Beltrami operator $-\Delta$ on $X$ is an increasing sequence
@@
It was remarked in [28, 15] that $-\log\mathrm{det}_{\zeta}\Delta$ can be interpreted as the total mass of the Brownian loop measure. The following expression from [1] gives a precise renormalization procedure, which shows that the zeta-regularized determinant of the Laplacian appears in the constant term in the truncation of the total mass of the Brownian loop measure after removing the small and large loops.
@@
And directly from this, Naud [37, Eq. (3)] shows that $-\log\mathrm{det}_{\zeta}\Delta$ can be written as the sum of length spectrum
$-\log\mathrm{det}_{\zeta}\Delta=-{\rm Area}(X)E-\upgamma+\int_{0}^{1}\frac{S_{X}(t)}{t}\,\mathrm{d}t+\int_{1}^{\infty}\frac{S_{X}(t)-1}{t}\,\mathrm{d}t$ (4.10)

### v3

@@ Abstract
We prove a simple identity relating the length spectrum of a Riemann surface to that of the same surface with an arbitrary number of additional cusps. Our proof uses the Brownian loop measure introduced by Lawler and Werner. In particular, we express the total mass of Brownian loops in a fixed free homotopy class on any Riemann surface in terms of the length of the geodesic representative for the complete constant curvature metric. This expression also allows us to write the electrical thickness of a compact set in $\mathbb{C}$ separating 0 and $\infty$, or the Velling–Kirillov Kähler potential, in terms of the Brownian loop measure and the zeta-regularized determinant of Laplacian as a renormalization of the Brownian loop measure with respect to the length spectrum.
@@ 1 Introduction
Let $(X,g)$ be a Riemannian surface without boundary (or we replace $X$ with boundary by $X\setminus\partial X$). From the uniformization theorem, if $X$ is not one of $\widehat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$, $\mathbb{C}$, $\mathbb{C}\setminus\{0\}$, and tori, then $X$ admits a unique complete hyperbolic metric that is conformal to $g$. In particular, $X$ may have cusps and funnels and may be of infinite type. We say a free homotopy class of oriented closed curves on $X$ is essential if it is not a trivial homotopy class nor a class homotopic to a cusp. Every such homotopy class of oriented closed curves on $X$ contains a unique hyperbolic geodesic $\gamma$. We define the iteration number $m(\gamma)\in\mathbb{Z}_{>0}$ if $\gamma$ is $m(\gamma)$-time iteration of a primitive closed geodesic.
@@ 1.2 Explicit computations of Brownian loop measure
Another immediate consequence is the following relation between the electrical thickness [22] and Brownian loop measure. Let $K\subset\mathbb{C}\setminus\{0\}$ be a connected compact set separating 0 and $\infty$. Let $\Omega$ and $\Omega^{*}$ be respectively the connected component of $\mathbb{C}\setminus K$ containing 0 and $\infty$. Let $f$ be a conformal map $\mathbb{D}\to\Omega$ fixing 0 and $h$ a conformal map $\mathbb{D}^{*}\to\Omega^{*}$ fixing $\infty$, where $\mathbb{D}=\{z\in\widehat{\mathbb{C}}\,|\,|z|<1\}$ and $\mathbb{D}^{*}=\{z\in\widehat{\mathbb{C}}\,|\,|z|>1\}$. The electrical thickness of $K$ is defined as
@@
Let $K\subset\mathbb{C}\setminus\{0\}$ be a connected compact set separating 0 and $\infty$. For all $m\in\mathbb{Z}\setminus\{0\}$, we have
@@
$\mu_{T}\{\text{loops homotopic to }\gamma^{m}\text{ and hit }\gamma\}=\frac{\operatorname{Area}(T)}{\pi\ell_{\gamma}(T)^{2}m^{2}}-\frac{1}{|m|}\frac{1}{\exp(\frac{\pi\ell_{\gamma}(T)^{2}|m|}{\operatorname{Area}(T)})-1},$
@@ 1.3 Zeta-regularized determinant of Laplacian
We show another way to express $\log\mathrm{det}_{\zeta}\Delta_{g}$ on a closed hyperbolic surface $X$ as a renormalization of the Brownian loop measure. The following result is a rewriting of the Selberg trace formula, which builds a connection between the Laplacian spectrum and the length spectrum, and hence is helpful to translate $\log\mathrm{det}_{\zeta}\Delta$ into the language of Brownian loop measure. Denote $\mathcal{G}(X)$ the set of all oriented closed geodesics on $X$, and $\mathcal{P}(X)\subset\mathcal{G}(X)$ the subset of primitive ones. Let $N_{X}(L):=\#\{\gamma\in\mathcal{P}(X)\mid\ell(\gamma)\leq L\}$.
@@
[…]
@@ 3 Brownian loop measure in a homotopy class
Since the Brownian loop measure is conformally invariant, choosing a complete metric with constant curvature on $X$ allows us to explicitly compute the total mass. In particular, $X$ is stochastically complete (i.e., Brownian motion is defined for infinite time, see Appendix A). If $X$ is not hyperbolic, then $X$ is conformal to one of $\widehat{\mathbb{C}},\mathbb{C},\mathbb{C}\setminus\{0\}$ or flat tori. On $\widehat{\mathbb{C}}$$\mathbb{C}$ and $\mathbb{C}\setminus\{0\}$, every homotopy class of closed curves is either homotopically trivial or homotopic to a puncture point, and hence the mass of Brownian loops is infinite. So, we will only consider complete hyperbolic surfaces and flat tori.
@@ 3.1 On hyperbolic surfaces
[…]
@@
where $\gamma]$ is the conjugacy class of $\gamma$ in $\Gamma$ and $z\in\mathbb{H}$ is any lift of $x$.
@@
[…]
@@
Fix an element $h_{0}\in[\gamma]$ and then $\gamma]=\{h_{1}^{-1}h_{0}h_{1}\mid h_{1}\in\Gamma\}$. Two elements $h_{1}^{-1}h_{0}h_{1}=h_{2}^{-1}h_{0}h_{2}$ if and only if $h_{2}h_{1}^{-1}h_{0}=h_{0}h_{2}h_{1}^{-1}$, that is, $h_{2}h_{1}^{-1}$ is in the centralizer of $h_{0}$ in $\Gamma$. Since $\Gamma$ is a discrete subgroup of ${\rm PSL}(2,\mathbb{R})$, the centralizer of $h_{0}\in\Gamma$ is just the set $\left\langle\delta\right\rangle:=\{\delta^{k}\mid k\in\mathbb{Z}\}$ where $\delta$ is primitive and $h_{0}=\delta^{m(\gamma)}$. Then
@@
where for each class $h_{1}]\in\left\langle\delta\right\rangle\backslash\Gamma$ we may choose any representation element.
@@
$U(u)$ $:=$ $\int_{\mathop{\rm arccosh}(u^{2})}^{\infty}\frac{r\mathrm{e}^{-r^{2}/4t}}{\sqrt{\cosh{r}-u^{2}}}\,\mathrm{d}r$ (3.3)
@@
$=\frac{\mathrm{e}^{-t/4}}{(4\pi t)^{3/2}}\frac{L}{m\sinh\frac{L}{2}}\ 2\int_{\sqrt{1+2(\sinh\frac{L}{2})^{2}}}^{\infty}\frac{U(u)u}{\sqrt{u^{2}-(1+2(\sinh\frac{L}{2})^{2})}}\,\mathrm{d}u$
@@ 4.2 On surfaces with boundary
As we mentioned at the beginning of Section 3, we may also use Lemma 3.2 to compute the Brownian loop measure on a surface $X$ with a non-empty boundary (where the Brownian loops are killed while hitting the boundary). Let $X^{\prime}=X\setminus\partial X$. Since $X^{\prime}$ cannot be one of $\widehat{\mathbb{C}},\mathbb{C},\mathbb{C}\setminus\{0\}$ or flat torus, it admits a unique complete hyperbolic metric. Then, by Theorem 2.2 and Lemma 3.2, the Brownian loop measure can be computed on $X^{\prime}$ with the complete hyperbolic metric:
@@
Let $K\subset\mathbb{D}$ be a closed subset such that $D=\mathbb{D}\setminus K$ is simply connected and contains 0. Let $\psi:D\to\mathbb{D}$ be a conformal map fixing 0. The total mass of Brownian loop in $\mathbb{D}$ intersecting $K$ with winding number $m$ around 0 equals $\log|\psi^{\prime}(0)|/(2\pi^{2}m^{2})$.
@@
Since $\psi$ is conformal, for all $\varepsilon>0$, there exists $r_{0}$ large enough, such that for all $r>r_{0}$, the curve $\psi(\mathrm{e}^{-r}\mathbb{S}^{1})$ is contained in the annulus $\{z\in\mathbb{C}\,|\,|\psi^{\prime}(0)|\mathrm{e}^{-r-\varepsilon}<|z|<|\psi^{\prime}(0)|\mathrm{e}^{-r+\varepsilon}\}$. See Figure 2 for an illustration. As $\mathbb{A}_{r^{\prime}}$ is conformally equivalent to the annulus bounded by $\mathbb{S}^{1}$ and $\psi(\mathrm{e}^{-r}\mathbb{S}^{1})$, we have from the monotonicity of the modulus of annuli for inclusion that $\log|\psi^{\prime}(0)|+r-\varepsilon<r^{\prime}<-\log|\psi^{\prime}(0)|+r+\varepsilon$. This implies $\lim_{r\to\infty}r-r^{\prime}=\log|\psi^{\prime}(0)|$.
@@
Let $K\subset\mathbb{C}\setminus\{0\}$ be a compact set separating 0 and $\infty$. Let $\Omega$ and $\Omega^{*}$ be respectively the connected component of $\mathbb{C}\setminus K$ containing 0 and $\infty$. Let $f$ be a conformal map $\mathbb{D}\to\Omega$ fixing 0 and $h$ a conformal map $\mathbb{D}^{*}\to\Omega^{*}$ fixing $\infty$. The electrical thickness [22] of $K$ (or the Velling–Kirillov Kähler potential in the context of universal Teichmüller theory [50]) is defined as
@@
Let $K\subset\mathbb{C}\setminus\{0\}$ be a connected compact set separating 0 and $\infty$. For all $m\in\mathbb{Z}\setminus\{0\}$, we have
@@
We write $\iota:z\mapsto 1/z$ and let $\tilde{h}=\iota\circ h\circ\iota$ which maps $\mathbb{D}$ conformally onto $\iota(\Omega^{*})$ while fixing 0. Then we have $|h^{\prime}(\infty)|=1/|\tilde{h}^{\prime}(0)|$ and $\theta(K)=-\log|\tilde{h}^{\prime}(0)|-\log|f^{\prime}(0)|$. The left-hand side of (4.6) equals
@@
[…]
@@
where $\delta>0$ is the Hausdorff dimension of the limit set of $\Gamma$. If $\delta>\frac{1}{2}$, then $\delta(1-\delta)$ is the smallest eigenvalue of $\Delta$ (acting on $L^{2}(X)$) on $X$ [40, 47, 48, 49]. In our case, $X$ has infinite area and hence constant valued functions are not $L^{2}$-eigenfunctions. So it has no zero eigenvalue (since the eigenfunction of zero eigenvalue is of constant value), and $\delta$ is strictly smaller than $1$. Then, the total mass of Brownian loops (4.7) is finite. ∎
@@ 4.3 Zeta-regularized determinant of Laplacian
Let us first recall briefly the definition of the zeta-regularized determinant. The spectrum of the positive Laplace–Beltrami operator $\Delta$ on $X$ is an increasing sequence
@@
It was remarked in [28, 15] that $\log\mathrm{det}_{\zeta}\Delta$ can be interpreted as the total mass of the Brownian loop measure. The following expression from [1] gives a precise renormalization procedure, which shows that the zeta-regularized determinant of the Laplacian appears in the constant term in the truncation of the total mass of the Brownian loop measure after removing the small and large loops.
@@
And directly from this, Naud [37, Eq. (3)] shows that $\log\mathrm{det}_{\zeta}\Delta$ can be written as the sum of length spectrum
$\log\mathrm{det}_{\zeta}\Delta=-{\rm Area}(X)E-\upgamma+\int_{0}^{1}\frac{S_{X}(t)}{t}\,\mathrm{d}t+\int_{1}^{\infty}\frac{S_{X}(t)-1}{t}\,\mathrm{d}t$ (4.10)